the rate of effusion of ethane gas is 1.9 times the rate of effusion of a partially decomposed Cl2O7 mixture .determine the degree of dissociation of Cl2O7?

Grade:12th pass

2 Answers

Arun
3 years ago
The relative rate of effusion is inversely proportional to the square root of molar mass.It is directly proportional to the number of moles .The molar masses of sulphur dioxide and methane are 64 g/mol and 16 g/mol respectively.
(i) Under similar condition of pressure & temperature
rCH4rSO2=MSO2MCH4=6416=21
( ii ) Through a container containing SO2 and CH4 in 3 : 2 mass ratio
rCH4rSO2=nCH4MSO2nSO2MCH4=162×64643×16=163
( iii ) If the mixture obtained by effusing out a mixture (nSO2/nCH4=8/1) for three effusing steps.
rCH4rSO2=MSO2MCH4=6416=
Vikas TU
3 years ago
Dear student 
Question is not clear 
Please attach an image, 
We will happy to  help you 
Good Luck
Cheers